### v1

@@ 5.3 Convergence, accuracy and computational cost of the LBFGS, and SL/TL-LBFGS optimizers
AC $3.6\times 10^{-3}$ $14{,}171{,}742$ $167{,}322$ $84.7$ $\boldsymbol{106{,}580}$ $132.9$ $1.6$
DA $5.5\times 10^{-1}$ $6{,}461{,}000$ $4{,}154$ $1555.4$ $\boldsymbol{533}$ $12,121.9$ $2.9$
Burg $1.5\times 10^{-4}$ $1{,}298{,}700$ $1{,}878{,}137$ $0.69$ $\boldsymbol{365{,}496}$ $3.5$ $5.1$
Aniso $1.7\times 10^{-2}$ $1{,}698{,}099$ $\boldsymbol{13{,}248}$ $128.2$ $51{,}168$ $33.2$ $0.3$
Adv $3.0\times 10^{-2}$ $472{,}394$ $16{,}072$ $29.4$ $\boldsymbol{14{,}130}$ $33.4$ $1.1$
Helm $7.3\times 10^{-2}$ $389{,}545$ $11{,}400$ $34.2$ $\boldsymbol{8{,}560}$ $45.5$ $1.3$
@@ Appendix A Condition number of (layer-wise) neural tangent kernel
Table 7 reports the condition number of neural tangent kernel (NTK) [12, 31], denoted by cond$(\mathbb{A})$ for all benchmark problems considered in Section 4. As we can see, the problems associated with training PINNs are more ill-conditioned than those associated with training DONs. Moreover, we also see that for PINN problems, the layer-wise condition number, denoted by $\text{cond}(\mathbb{A}_{l})$, varies over a wider range. This suggests that the local training associated with solving (11) requires different adjustments of step-size. Thus, by decoupling the layers, we are able to use larger step-sizes for layers that are better conditioned, which in turn enhances the overall convergence.
@@
Helm $21$ $18$ $21$

### v2

@@ 5.3 Convergence, accuracy and computational cost of LBFGS and SL/TL-LBFGS optimizers
AC $3.6\times 10^{-3}$ 14,171,74214{,}171{,}742 167,322167{,}322 $84.7$ $\boldsymbol{106{,}580}$ $132.9$ $1.6$
DA $5.5\times 10^{-1}$ 6,461,0006{,}461{,}000 4,1544{,}154 $1555.4$ $\boldsymbol{533}$ $12,121.9$ $2.9$
Burg $1.5\times 10^{-4}$ 1,298,7001{,}298{,}700 1,878,1371{,}878{,}137 $0.69$ $\boldsymbol{365{,}496}$ $3.5$ $5.1$
Aniso $1.7\times 10^{-2}$ 1,698,0991{,}698{,}099 $\boldsymbol{13{,}248}$ $128.2$ 51,16851{,}168 $33.2$ $0.3$
Adv $3.0\times 10^{-2}$ 472,394472{,}394 16,07216{,}072 $29.4$ $\boldsymbol{14{,}130}$ $33.4$ $1.1$
Helm $7.3\times 10^{-2}$ 389,545389{,}545 11,40011{,}400 $34.2$ $\boldsymbol{8{,}560}$ $45.5$ $1.3$
6 Discussion
Machine learning (ML)-based solvers for PDEs are rapidly gaining attention; however, their role remains distinct from that of classical numerical methods. Established approaches such as finite element and finite volume methods, when preconditioned by multigrid or domain decomposition techniques, continue to be the most efficient and reliable tools for solving large-scale, three-dimensional problems on unstructured geometries. At present, machine learning-based solvers should be regarded primarily as complementary methods, with particular strengths in high-dimensional problems, inverse problems, model discovery, and surrogate modeling for parametric problems.
A major drawback of current machine learning-based approaches is the absence of rigorous error analysis and convergence guarantees, which makes it challenging to assess their reliability compared to classical methods. Their performance is also highly sensitive to heuristic design choices, including the architecture of the network, the selection of hyperparameters, and the training strategy. From a computational standpoint, classical methods can often solve benchmark problems in fractions of a second, whereas training machine learning-based solvers typically require significant time and computational resources. As demonstrated in the manuscript, the development of novel, large-scale, training strategies can significantly reduce the associated computational burden. Moreover, once trained, neural networks enable extremely fast evaluations across multiple parameter configurations, which is particularly appealing for applications such as shape optimization, design under uncertainty, or real-time simulation.
In summary, although machine learning-based solvers are not yet ready to fully replace state-of-the-art numerical methods, they represent a promising complementary paradigm. A key open challenge is the development of novel training strategies that combine improved efficiency with enhanced model accuracy, ultimately enabling machine learning-based solvers to evolve into practical and reliable tools for large-scale scientific computing.
@@ 7 Conclusion
Acknowledgements
Y.L. was supported in part by Basic Science Research Program through NRF funded by the Ministry of Education (No. RS2023-00247199). A.K. was supported by the Swiss National Science Foundation (SNSF) under the project “Multilevel training of DeepONets – multiscale and multiphysics applications”, as well as by the Platform for Advanced Scientific Computing (PASC) under the project ExaTrain. Moreover, work of A.K. benefited from the AI Interdisciplinary Institute ANITI, funded by the France 2030 program under Grant Agreement No. ANR-23-IACL-0002. G.E.K. is supported by the ONR Vannevar Bush Faculty Fellowship. We also acknowledge support from the DOE-MMICS SEA-CROGS DE-SC0023191 award and Ansys Inc.
@@ Appendix A Condition number of (layer-wise) neural tangent kernel
[…]
@@
Helm 2121 $18$ 2121
Appendix B Getting started with DistTrainNN
In this section, we briefly describe how to run an TL-LBFGS training algorithms within our open-source  DistTraiNN [66]. Firstly, the user is required to setup a Python environment with the following depedencies: PyTorch, Pandas, pyDOE, scikit-optimize, scikit-learn.
The scripts used to obtain the numerical results reported in Section 5 are available in the folder “tests/SPQN_tests”. An example code for training a PINN to approximate a solution for the Allen–Cahn equation is provided below. Please, note that the optimizer is fully configurable via command line arguments.
git clone git@bitbucket.org:alena_kopanicakova/disttrainn.git
cd disttrainn
export PYTHONPATH="\${PYTHONPATH}:/path/to/disttrainn"
cd tests/TL_SPQN_tests/AllenCahn
python3 test_SPQN_twolevel_additive_serial.py --seed 1 --num_points_x 25 --num_points_y 50 --num_ref 4 --max_epochs 200000 --lr_global 1.0 --lr_local 1.0 --history 3 --width 64 --hiddenlayers_coarse 4 --num_subdomains 3 --num_local_steps 50 --num_coarse_steps 50 --use_layerwise_lr True --overlap_width 1